The mass of the sun is 1.991x10^30 kg and its density is 1.410 g/cm^3. What is the volume of the sun in m^3?
1
Expert's answer
2013-06-27T07:39:35-0400
Mass, volume and density of smth are related by the following formula:
Mass=Volume*Density
Therefore, Volume=Mass/Density
Also, we need to convert given density into kg/m^3:
1.410 g/cm^3=1.410*10^(-3)/10^(-6) kg/m^3=1410 kg/m^3

So in our case we have Volume=1.991*10^30 kg/1410 kg/m^3=1412*10^(24) m^3